Dustylove2020:

Paul runs 4.8 km every morning. How many feet are in 4.8 km, given that 1 mile = 1.609 km and 1 mile = 5,280 feet?

Mercury:

\[\frac{ 4.8km }{ }\frac{ 1~mile }{ 1.609km }\frac{ 5,280~feet }{ 1~mile }\] notice how I've set up the conversion factors such that units cancel out, and you're left with feet